solve the following for x by any method find all solutions, real or complex. write your answer in simplest radical form, and in terms of i 6(x-5)(x-5)=42
Answer by richwmiller(17219)

(x-5)^2=42/6=7
x-5=sqrt(7)
x=5+sqrt(7)
x=5-sqrt(7)